# Chance-Constrained DC Optimal Power Flow Using Constraint-Informed Statistical Estimation

### v1

@@ Abstract
Index Terms:
Chance Constraints, DC Optimal Power Flow
@@ I Introduction
However, several statistical analyses have shown that non-Gaussian distributions, such as Weibull [15], [16], Cauchy [17] (and Cauchy-like non-Gaussian distribution errors [18], [19]), may be better capable of representing wind generation power forecast errors, particularly within the shorter time horizon of OPF applications. Unfortunately, from the optimization perspective, the incorporation of non-Gaussian distributions is not as computationally straightforward as in the Gaussian case, since the resulting CC-OPF problems do not admit tractable reformulations that can be solved by existing solvers.
A promising path for modeling these non-Gaussian forecast errors is via Gaussian Mixture Models (GMMs), as they can approximate a wide class of probability distributions by mixing a sufficiently large number of Gaussian components [20]. At the same time, the analytical reformulation technique for Gaussian chance constraints has also been recently extended to the GMM setting [21, 22, 23, 24, 25, 26], enabling solution of these models using standard solvers.
However, a key drawback of these and other existing approaches is that they use a multi-dimensional GMM to model forecasting errors. In particular, the dimension of the forecasting error distribution scales with the number of buses with stochastic injections (e.g., wind farms). From a statistical standpoint, this approach presents several challenges. First, standard MLE algorithms become significantly harder for high-dimensional GMMs. Unlike in the case of Gaussians, the MLE objective function is now nonconvex, and the Expectation-Maximization (EM) algorithm typically used in this context is only heuristic and prone to converge to local solutions [27]. Second, the problem is affected by the curse of dimensionality: the number of parameters to be estimated can grow so large that historical data becomes insufficient, resulting in models that are prone to overfitting and that fail to generalize to unseen future data [28, 29]. In particular, roughly each additional $n$-dimensional Gaussian component in a GMM requires $O(n^{2})$ new parameters to be estimated.
@@
Building on the constraint-informed framework, we also develop tractable analytical reformulations of chance constraints when the low-dimensional uncertainties follow a GMM. Through case studies, we show that it is the combination of constraint-informed dimensionality reduction and Gaussian mixture modeling that delivers significantly improved out-of-sample reliability compared to existing approahces. We find that the improvement is particularly stark when the underlying forecast errors exhibit unimodal but heavy-tailed behavior, elucidating that GMMs can also reduce out-of-sample risk in systems that may not involve multimodal data.
@@ II Chance-Constrained DC-OPF Model
Note that if we assume $\xi_{i}$ are i.i.d. zero-mean Gaussian random variables with known variance, then (16) further simplifies and reduces to the objective function used in prior work [6, 12]. Constraints (15b) and (15c) together ensure power balance within the network under the affine control policy. Constraints (15d) and (15e) are chance constraints enforcing that the power generated lies within the minimum and maximum generation limits with (sufficiently high) probability $1-\epsilon$. Similarly, constraints (15f) and (15g) are chance constraints enforcing power flows along lines to remain less than the maximum line capacity with probability $1-\epsilon$. Here, $\epsilon$ is a risk parameter to be determined by the system operator. Lower values of $\epsilon$ reflect higher risk aversion, since they enforce a higher probability of constraint satisfaction. We refer the reader to related works [35, 13] for setting $\epsilon$ values.
@@ III Constraint-Informed Reformulation
For brevity, we present our ideas in the context of constraints (15e) and (15g) only. The corresponding techniques for (15d) and (15f) are entirely analogous.
@@ III-B Constraint-Informed Approach: Transform and then Fit
    Use the fitted parameters to to obtain the distributions of the random variables, $p_{g}(\bm{\xi})$ and $f_{l}(\bm{\xi})$, and to reformulate the chance constraints (15e) and (15g), respectively:
@@
When making dispatch decisions, system operators are primarily concerned with system-wide imbalances rather than individual unit deviations, thus directly modeling $\Omega$ is also operationally meaningful. On the other hand, the construction of data samples for $\bm{\eta}_{l}$ from raw forecast error data takes into consideration both aggregate system-wide forecast errors $\Omega$ and line-specific localized forecast errors $\Lambda_{l}$. Therefore, our constraint-informed approach efficiently captures both the global and local effects of forecast uncertainty, while avoiding high-dimensional statistical estimation. Fig. 1 shows the flowchart of the classical and constraint-informed approaches.
Original Data Samples $\{\bm{\xi}^{(n)}\}_{n=1}^{N}$High-Dimensional Statistical Fitting $\hat{\pi}_{k},\hat{\bm{\mu}}_{k},\hat{\bm{\Sigma}}_{k}$Transformed Parameters in One Dimension $\bm{1}^{\top}\hat{\bm{\mu}}_{k},\bm{1}^{\top}\hat{\bm{\Sigma}}_{k}\bm{1}$Reformulated Chance ConstraintsPower Generation DecisionsLow-Dimensional Data Sample Construction $\{\Omega^{(n)}\}_{n=1}^{N},\{\bm{\eta}_{l}^{(n)}\}_{n=1}^{N}$Low-Dimensional Statistical Fitting $\hat{\beta}_{k},\hat{m}_{k},\hat{\sigma}_{k}^{2},\hat{\lambda}_{k},\hat{\bm{\nu}}_{k},\hat{\bm{C}}_{k}$Reformulated Chance ConstraintsPower Generation Decisions Classical Approach Constraint-Informed Approach
Figure 1: Comparison of Classical and Constraint-Informed Approaches for Reformulating Chance Constraints
@@ III-E GMM Chance Constraint Reformulation
To address this challenge, we adapt an existing method from the literature [25] that builds a piecewise linear (PWL) approximation of the standard normal CDF $\Phi$. The key idea is to discretize $\Phi$ over the nonnegative reals into $S$ linear segments, using breakpoints $t_{0}=0<t_{1}<\cdots<t_{S-1}$. These breakpoints can be chosen optimally based on a user-specified tolerance $\delta$ [25, Algorithm 1] so that the linear approximation $\hat{\Phi}$ remains as close as possible to $\Phi$; namely, $\lVert\hat{\Phi}(x)-\Phi(x)\rVert\leq\delta,\forall x\in[0,\infty)$, with a minimal number of segments. For each segment defined over the interval $t_{s-1},t_{s}]$, we approximate $\Phi$ from below with a linear function whose slope $a_{s}$ and intercept $b_{s}$ can be calculated as follows:
@@
The pointwise minimum ensures that the PWL approximation $\hat{\Phi}$ is a valid under-estimator for $\Phi$, so that the solution obtained by substituting $\Phi$ with $\hat{\Phi}$ in constraint (25) will also be feasible for the original constraint (15e).
@@
We can multiply both sides by (the nonnegative term) $\alpha_{g}$ to eliminate nonlinearity in the constraint. Similarly, we can introduce an auxiliary variable $M_{gk}^{1}$ for each piecewise linear segment and each $g\in\mathcal{G}$ to eliminate the minimum operator.
@@
Convexity of the reformulation, however, requires the additional assumption that the component-specific covariance matrices of $\bm{\eta}_{l}$ share the same covariance matrix [^2] This is supported by standard statistical software; e.g., spherical or tied covariance in the sklearn.mixture Python package.: $\hat{\bm{C}}_{k}=\tau_{k}^{2}\hat{\bm{C}}_{0}$. Utilizing techniques similar to (34) and (35), we obtain:
@@ IV-A Test Systems and Datasets
We perform experiments on a modified IEEE 118 bus test case. In particular, we replace 10 out of 19 conventional generators with wind units that don’t incur any generation cost. We set the chance constraint risk threshold to $\epsilon=0.05$. The accuracy of the piecewise linear approximation of $\Phi$ is set to $0.002$, which results in $S=10$ segments being used.
We experiment with three datasets $\{\bm{\xi}^{(n)}\}_{n=1}^{N}$ of forecast errors: synthetic Gaussian-distributed errors (Synthetic-G), synthetic Cauchy-distributed errors (Synthetic-C), and real-world NordPool data [39]. For the synthetic cases, we generate ten independent datasets of forecast errors using distributional parameters [^3] Gaussian: $\mu=-0.024,\sigma=0.036$ and Cauchy: $x_{0}=0,\gamma=0.02$. adapted from existing literature [26, 17]. Each dataset consists of 10,000 samples, with $N=8,000$ samples used for statistical estimation and subsequent optimization, while the remaining 2,000 for evaluating out-of-sample risk.
@@ IV-B Goodness-of-Fit
For numerical comparisons, we compute the log-likelihood of the two densities to quantify the similarity of the estimated distributions to the empirical aggregate forecast errors. A higher log-likelihood indicates a better fit of the distribution to the observed data. Table III reports the best log-likelihood values, obtained from ten different initializations of the EM algorithm across ten different splits of each dataset. The corresponding density curves are plotted in Figure 3.
TABLE III: Log-likelihood of $\Omega$ (higher is better)
Dataset, Fitted Model Classical Constraint-Informed
Synthetic-G, GMM ($K=1$) 7367 7367
Synthetic-C, GMM ($K=3$) -33649 -9869
NordPool, GMM ($K=3$) -4739 -3343
In the Synthetic-G dataset, where forecast errors are normally distributed, we find that the classical and constraint-informed approaches yield equal log-likelihood values and identical overlapping density curves, validating the MLE argument from Section III-C. Unsurprisingly, the positive log-likelihood value indicates a good Gaussian fit in this case.
@@
The results for the Synthetic-C dataset highlight the limitations of the classical approach. We believe this can be explained as follows: since the Cauchy distribution has undefined first and second moments, the conditions for the Central Limit Theorem do not apply. As a result, summing Cauchy-distributed errors does not approximate a normal distribution. The classical method, which fits a high-dimensional distribution and then reduces it to one dimension, overemphasizes the heavy tails by estimating a high-variance Gaussian component. It is possible to improve the log-likelihood using a modified EM algorithm [36]; we present this detail in Appendix C.
The failure of the classical approach is evidenced in the middle and right-hand side plots of Figure 3, where the density curves deviate significantly from the true histograms of the Synthetic-C and NordPool datasets. In contrast, the constraint-informed approach, which directly fits a distribution to one-dimensional aggregated data, provides a much better match to the true samples. This is also supported by the log-likelihood values in Table III.

### v2

@@ I Introduction
However, several statistical analyses have shown that non-Gaussian distributions, such as Weibull [15], Cauchy [16], generalized Pareto [17], (and Cauchy-like non-Gaussian distribution errors [18], [19]), may be better capable of representing wind generation power forecast errors, particularly within the shorter time horizon of OPF applications. Unfortunately, from the optimization perspective, the incorporation of non-Gaussian distributions is not as computationally straightforward as in the Gaussian case, since the resulting CC-OPF problems do not admit tractable reformulations that can be solved by existing solvers.
A promising path for modeling these non-Gaussian forecast errors is via Gaussian Mixture Models (GMMs), as they can approximate a wide class of probability distributions by mixing a sufficiently large number of Gaussian components [20]. At the same time, the analytical reformulation technique for Gaussian chance constraints has also been recently extended to the GMM setting [21, 22, 23, 24, 25, 26], enabling solution of these models using standard solvers. The key advantage of GMMs is that linear combinations of GMM-distributed random variables remain GMM-distributed, allowing chance constraints to be reformulated analytically using component-specific Gaussian parameters. In contrast, non-GMM distributions such as Weibull, Cauchy, or generalized Pareto, typically do not preserve their distributional form under linear combinations (e.g., the sum of two Weibull random variables is not Weibull), precluding direct analytical reformulation. Therefore, we use GMMs to approximate forecast errors, enabling both analytical reformulations and the ability to capture non-Gaussian behavior.
In this vein, [21] fits a GMM to the joint probability density of wind power output across multiple farms for chance-constrained economic dispatch; [23] uses a customized GMM to approximate the probability density of wind power generation for probabilistic OPF; [24] uses GMMs as input distributions in stochastic power flow and state estimation of distribution networks; [22] uses a generalized GMM to characterize wind power ramp distributions; and [25] and [26] fit GMMs to wind power generation at individual units for CC-OPF reformulation. In all these approaches, GMMs are fitted to the raw uncertainty at the individual component level, and the resulting distribution parameters are subsequently transformed to evaluate and reformulate the chance constraints.
However, a key drawback of this paradigm is that the joint distribution of component-level uncertainties is high-dimensional. In particular, the dimension of the forecasting error distribution scales with the number of buses with stochastic injections (e.g., wind farms). From a statistical standpoint, this approach presents several challenges. First, standard MLE algorithms become significantly harder for high-dimensional GMMs. Unlike in the case of Gaussians, the MLE objective function is now nonconvex, and the Expectation-Maximization (EM) algorithm typically used in this context is only heuristic and prone to converge to local solutions [27]. Second, the problem is affected by the curse of dimensionality: the number of parameters to be estimated can grow so large that historical data becomes insufficient, resulting in models that are prone to overfitting and that fail to generalize to unseen future data [28, 29]. In particular, roughly each additional $n$-dimensional Gaussian component in a GMM requires $O(n^{2})$ new parameters to be estimated.
@@
Building on the constraint-informed framework, we also develop tractable analytical reformulations of chance constraints when the low-dimensional uncertainties follow a GMM. Through case studies, we show that it is the combination of constraint-informed dimensionality reduction and Gaussian mixture modeling that delivers significantly improved out-of-sample reliability compared to existing approaches. We find that the improvement is particularly stark when the underlying forecast errors exhibit unimodal but heavy-tailed behavior, elucidating that GMMs can also reduce out-of-sample risk in systems that may not involve multimodal data.
@@ II Chance-Constrained DC-OPF Model
Note that if we assume $\xi_{i}$ are i.i.d. zero-mean Gaussian random variables with known variance, then (16) further simplifies and reduces to the objective function used in prior work [6, 12]. Constraints (15b) and (15c) together ensure power balance within the network under the affine control policy. Constraints (15d) and (15e) are chance constraints enforcing that the power generated lies within the minimum and maximum generation limits with (sufficiently high) probability $1-\epsilon$. Similarly, constraints (15f) and (15g) are chance constraints enforcing power flows along lines to remain less than the maximum line capacity with probability $1-\epsilon$. Here, $\epsilon$ is a risk parameter to be determined by the system operator. Lower values of $\epsilon$ reflect higher risk aversion, since they enforce a higher probability of constraint satisfaction. In practice, different $\epsilon$ values may be assigned to different classes of constraints to reflect their operational importance. We refer the reader to related works [34, 12, 35, 13] for further discussion on setting $\epsilon$ values; in our experiments, we use a common value across all constraints so that the comparison focuses on the estimation methodology.
@@ III Constraint-Informed Reformulation
Let $\bm{z}=(\bar{\bm{p}},\bm{\alpha})$ denote the decision variables of (15). Each chance constraint in (15d)–(15g) can be written in the form
$\mathbb{P}\big(\bm{a}(\bm{z})^{\top}\bm{A}\bm{\xi}\leq b(\bm{z})\big)\geq 1-\epsilon,$ (17)
where $\bm{a}(\bm{z})$ and $b(\bm{z})$ are affine functions of $\bm{z}$ and $\bm{A}$ is a constant matrix. For the power generation constraints (15e)–(15d), this follows from (1) and (2) with $\bm{A}=\bm{1}^{\top}\in\mathbb{R}^{1\times|\mathcal{B}|}$, so that $\bm{A}\bm{\xi}=\Omega\in\mathbb{R}$. For the line flow constraints (15g)–(15f), this follows from (8), (13), and (14) with $\bm{A}=\begin{bmatrix}\bm{1}&\bm{h}_{l}^{\mathrm{wind}}\end{bmatrix}^{\top}\in\mathbb{R}^{2\times|\mathcal{B}|},$ so that $\bm{A}\bm{\xi}=\bm{\eta}_{l}\in\mathbb{R}^{2}$.
The key observation is that the constraint (17) depends on uncertainty only through the projection $\bm{A}\bm{\xi}$, which is a low-dimensional linear function of the raw uncertainty $\bm{\xi}$. To evaluate the probability, one only needs the distribution of $\bm{A}\bm{\xi}$ and not of $\bm{\xi}$ itself. Since $\bm{A}\bm{\xi}$ is one-dimensional for power generation and two-dimensional for line flow constraints, learning its distribution is significantly more efficient and accurate than learning the full joint distribution of $\bm{\xi}\in\mathbb{R}^{|\mathcal{B}|}$.
We refer to this as the constraint-informed approach because the projection $\bm{A}$ is informed directly from the structure of the chance constraints. The matrix $\bm{A}$ characterizes precisely how the raw uncertainty $\bm{\xi}$ enters and influences each constraint. The approach applies to any linear chance-constrained model: one needs only to identify the low-dimensional projection $\bm{A}\bm{\xi}$ and fit a GMM distribution to the transformations $\{\bm{A}\bm{\xi}^{(n)}\}_{n=1}^{N}$ of the raw samples $\{\bm{\xi}^{(n)}\}_{n=1}^{N}$. Because the distribution is fit to low-dimensional data, it is more accurate and efficient to learn compared to fitting $\bm{\xi}$ directly and then transforming its fitted parameters. We accordingly refer to the constraint-informed approach as transform and then fit, in contrast to the classical fit and then transform. Fig. 1 illustrates this comparison.
For brevity, we present the detailed reformulation steps for the upper-bound constraints (15e) and (15g) only. The treatment of (15d) and (15f) is entirely analogous.
Raw Forecast Error Data $\{\bm{\xi}^{(n)}\}_{n=1}^{N}$Fit high-dimensional Gaussian mixture distribution to $\bm{\xi}$Transform fitted parameters to obtain distribution of $\bm{A}\bm{\xi}$Reformulated Chance ConstraintsOptimal Dispatch DecisionsTransform data samples: $\{\bm{A}\bm{\xi}^{(n)}\}_{n=1}^{N}$Fit low-dimensional Gaussian mixture distribution to $\bm{A}\bm{\xi}$Reformulated Chance ConstraintsOptimal Dispatch Decisions Classical ApproachFit and then Transform Constraint-Informed ApproachTransform and then Fit
Figure 1: Comparison of the classical (fit and then transform) and constraint-informed (transform and then fit) approaches for reformulating chance constraints. The matrix $\bm{A}$ denotes the low-dimensional linear projection identified from the constraint structure.
@@ III-B Constraint-Informed Approach: Transform and then Fit
    Use the fitted parameters to obtain the distributions of the random variables, $p_{g}(\bm{\xi})$ and $f_{l}(\bm{\xi})$, and to reformulate the chance constraints (15e) and (15g), respectively:
@@
When making dispatch decisions, system operators are primarily concerned with system-wide imbalances rather than individual unit deviations, thus directly modeling $\Omega$ is also operationally meaningful. On the other hand, the construction of data samples for $\bm{\eta}_{l}$ from raw forecast error data takes into consideration both aggregate system-wide forecast errors $\Omega$ and line-specific localized forecast errors $\Lambda_{l}$. Therefore, our constraint-informed approach efficiently captures both the global and local effects of forecast uncertainty, while avoiding high-dimensional statistical estimation. .
@@ III-E GMM Chance Constraint Reformulation
To address this challenge, we adapt an existing method from the literature [25] that builds a piecewise linear (PWL) approximation of the standard normal CDF $\Phi$. The key idea is to discretize $\Phi$ over the nonnegative reals into $S$ linear segments, using breakpoints $t_{0}=0<t_{1}<\cdots<t_{S-1}$. Given a user-specified tolerance $\delta>0$, these breakpoints are chosen optimally by [25, Algorithm 1] to minimize the number of segments $S$ while ensuring that $\lVert\hat{\Phi}(x)-\Phi(x)\rVert\leq\delta$ for all $x\in[0,\infty)$. Thus, $S$ is not a free parameter but is determined by the choice of $\delta$; a smaller $\delta$ yields a tighter approximation at the cost of more segments. We discuss how to set $\delta$ in Section IV. For each segment defined over the interval $t_{s-1},t_{s}]$, we approximate $\Phi$ from below with a linear function whose slope $a_{s}$ and intercept $b_{s}$ can be calculated as follows:
@@
The pointwise minimum ensures that the PWL approximation $\hat{\Phi}$ is a valid under-estimator for $\Phi$, so that the solution obtained by substituting $\Phi$ with $\hat{\Phi}$ in constraint (26) will also be feasible for the original constraint (15e). In other words, the PWL approximation always yields conservative (i.e., more reliable) solutions, regardless of the number of segments $S$.
@@
We can multiply both sides by (the nonnegative term) $\alpha_{g}$ to eliminate nonlinearity in the constraint. Similarly, we can introduce a continuous auxiliary variable $M_{gk}^{1}$ for each component $k\in[K]$ and each $g\in\mathcal{G}$ to eliminate the minimum operator; we emphasize that no binary variables are needed.
@@
Convexity of the reformulation, however, requires the additional assumption that the component-specific covariance matrices of $\bm{\eta}_{l}$ share the same covariance matrix [^2] This is supported by standard statistical software; e.g., spherical or tied covariance in the sklearn.mixture Python package. In our experiments, the tied structure achieved nearly identical log-likelihoods to unrestricted covariances while using fewer parameters, and this advantage is particularly pronounced when data are scarce.: $\hat{\bm{C}}_{k}=\tau_{k}^{2}\hat{\bm{C}}_{0}$. Relaxing this assumption for $\bm{\eta}_{l}$ would require introducing binary variables to handle arbitrary component covariances [37], destroying convexity and real-time applicability of the method. Utilizing techniques similar to (35) and (36), we obtain:
@@ IV-A Test Systems and Datasets
We perform experiments on a modified IEEE 118 bus test case. In particular, we replace 10 out of 19 conventional generators with wind units that don’t incur any generation cost. To assess scalability, we also consider a larger 2746-bus Polish system in Section IV-E. We set the chance constraint risk threshold to $\epsilon=0.05$. The accuracy of the piecewise linear approximation of $\Phi$ is set to $\delta=0.1\epsilon$. [^3] This choice can be explained as follows. Since $0\leq\Phi(x)-\hat{\Phi}(x)\leq\delta$ for all $x\geq 0$, enforcing the chance constraints with $\hat{\Phi}$ in place of $\Phi$ can be shown to introduce at most $\delta$ additional conservatism compared to the original constraint (since the mixture weights sum to one). The factor of $0.1$ ensures that this conservatism is at least one order of magnitude smaller than $\epsilon$. With $\epsilon=0.05$, this gives $\delta=0.005$, resulting in $S=6$ PWL segments [^4] Reducing $\delta$ results in more PWL segments $S$ that increase computational time but also provide a tighter approximation of the Gaussian CDF and hence a decrease in cost and solution conservatism. We find that tightening the tolerance to $\delta=0.002$ ($S=10$) and $\delta=0.001$ ($S=13$) increased solution time without any meaningful change in out-of-sample risk..
We experiment with three datasets $\{\bm{\xi}^{(n)}\}_{n=1}^{N}$ of forecast errors: synthetic Gaussian-distributed errors (Synthetic-G), synthetic Cauchy-distributed errors (Synthetic-C), and real-world NordPool data [39]. The choice of the Cauchy distribution is illustrative; the proposed framework applies to any continuous distribution since GMMs are universal approximators [20]. The forecast errors themselves need not follow a GMM and the latter is only used as a statistical approximation of the empirical data. For the synthetic cases, we generate ten independent datasets of forecast errors using distributional parameters [^5] Gaussian: $\mu=-0.024,\sigma=0.036$ and Cauchy: $x_{0}=0,\gamma=0.02$. adapted from existing literature [26, 16]. Each dataset consists of 10,000 samples, with $N=8,000$ samples used for statistical estimation and subsequent optimization, while the remaining 2,000 for evaluating out-of-sample risk.
@@ IV-B Goodness-of-Fit
For numerical comparisons, we compute the Bayesian Information Criterion (BIC) scores of the GMMs estimated using both approaches for $K=1,2,\dots,10$ Gaussian components. For each $K$, we estimate the GMM using 10 different EM initializations and record the lowest BIC. The BIC balances goodness-of-fit with model complexity: a lower BIC indicates a better fit relative to model size [36]. Table III reports the lowest BIC and associated optimal $K$ for each dataset and approach. The corresponding density curves are plotted in Figure 3.
TABLE III: Best BIC scores and optimal number of components $K$ for the aggregate forecast error $\Omega$ (lower BIC is better)
Dataset Classical Constraint-Informed
Synthetic-G 14761.09(K=1)-14761.09\ (K=1) 14716.09(K=1)-14716.09\ (K=1)
Synthetic-C 44009.11(K=9)44009.11\ (K=9) 8986.49(K=4)8986.49\ (K=4)
NordPool 5721.74(K=9)5721.74\ (K=9) 5517.40(K=6)5517.40\ (K=6)
In the Synthetic-G dataset, where forecast errors are normally distributed, we find that the classical and constraint-informed approaches select $K=1$ with nearly identical BIC values and overlapping density curves, validating the MLE equivalence argument from Section III-C.
@@
The results for the Synthetic-C dataset highlight the limitations of the classical approach. We believe this can be explained as follows: since the Cauchy distribution has undefined first and second moments, the conditions for the Central Limit Theorem do not apply. As a result, summing Cauchy-distributed errors does not approximate a normal distribution. The classical method, which fits a high-dimensional distribution and then reduces it to one dimension, overemphasizes the heavy tails by estimating a high-variance Gaussian component.
The failure of the classical approach is evidenced in the middle and right-hand side plots of Figure 3: even with more Gaussian components, the classical approach achieves substantially higher BIC scores than the constraint-informed approach, and its density curves deviate significantly from the true histograms of the Synthetic-C and NordPool datasets. In contrast, the constraint-informed approach, which directly fits a distribution to one-dimensional aggregated data, provides a much better match to the true samples. This is also supported by the BIC scores in Table III. This improvement in fitting is also achieved with a significantly smaller number of mixture components. For the Synthetic-C dataset, the constraint-informed approach requires only 4 Gaussian components, whereas the classical approach requires 9. Previous work [21] reported 7–9 Gaussian components for a GMM fit of $\bm{\xi}$ in a 39-bus system; our results both validate this observation and demonstrate that the proposed approach can achieve better accuracy with fewer components.